OpenStudy (anonymous):

Find the arc length of y=x^/16+1/2x^2 x=2 x=3

OpenStudy (anonymous):

Use this site I'll help https://www.wolframalpha.com/

OpenStudy (anonymous):

Thanks but I kinda want a more in depth explanation do how to do it

ganeshie8 (ganeshie8):

find the derivative, and plug it in arclength formula

OpenStudy (anonymous):

What would the derivative be? @ganeshie8

ganeshie8 (ganeshie8):

dy/dx = ?

ganeshie8 (ganeshie8):

you need help in finding the derivative ?

OpenStudy (anonymous):

Well when I took the derivative I got x^3/64-1/2x^2

OpenStudy (anonymous):

I don't think that's right

ganeshie8 (ganeshie8):

whats ur function ?

ganeshie8 (ganeshie8):

can u type it agian here ?

ganeshie8 (ganeshie8):

there seem to be some formatting error in ur original question

ganeshie8 (ganeshie8):

y=x^/16+1/2x^2

ganeshie8 (ganeshie8):

??

OpenStudy (anonymous):

Oh sorry it's y=x^4/16+1/2x^2

ganeshie8 (ganeshie8):

ty :) \(\large y = \frac{x^4}{16} + \frac{1}{2x^2}\)

ganeshie8 (ganeshie8):

like this ?

OpenStudy (anonymous):

Yep

ganeshie8 (ganeshie8):

\(\large y = \frac{x^4}{16} + \frac{1}{2x^2}\) \(\large \frac{dy}{dx} = \frac{d}{dx} \left(\frac{x^4}{16} + \frac{1}{2x^2}\right)\) \(\large ~~~~~= \frac{d}{dx} \left(\frac{x^4}{16}\right) + \frac{d}{dx}\left(\frac{1}{2x^2}\right)\) \(\large ~~~~~= \frac{1}{16}\frac{d}{dx} \left(x^4\right) + \frac{1}{2}\frac{d}{dx}\left(x^{-2}\right)\)

ganeshie8 (ganeshie8):

whats the derivative of x^4 ? whats the derivative of x^(-2) ?

OpenStudy (anonymous):

The derivative of x^4 is 4x^3

OpenStudy (anonymous):

And x^-2 is -2x^-3

OpenStudy (phi):

yes.

OpenStudy (phi):

now find dy/dx

OpenStudy (anonymous):

Uhh would it be 3/4x^2+3x^-4?

OpenStudy (phi):

it's \[ \frac{1}{16}\frac{d}{dx} \left(x^4\right) + \frac{1}{2}\frac{d}{dx}\left(x^{-2}\right) \] replace d/dx(x^4) with your result. same for the other

OpenStudy (anonymous):

Yeah I did that .-.

OpenStudy (phi):

you should get this \[ \frac{1}{16}\left(4 x^3\right) + \frac{-2}{2} \left(x^{-3}\right) \] which can be simplified

OpenStudy (anonymous):

1/4x^3+ (-1x^-3)

OpenStudy (phi):

yes. \[ y' = \frac{x^3}{4} - x^{-3} \] you now need (y')^2

OpenStudy (phi):

we are going to use the arc length formula see http://tutorial.math.lamar.edu/Classes/CalcII/ArcLength.aspx for info.

OpenStudy (anonymous):

Can I just distribute the 2 or do I have to foil it?

OpenStudy (phi):

you have to foil. (y')^2 means y' * y' or in this case \[ \left( \frac{x^3}{4} - x^{-3} \right)\left( \frac{x^3}{4} - x^{-3} \right)\]

OpenStudy (phi):

It looks a bit ugly, but we can deal with it.

OpenStudy (anonymous):

Okay let me see what I get

OpenStudy (anonymous):

X^6/16-1/2+x^-6?

OpenStudy (phi):

yes. the arc length formula uses \[ \sqrt{ 1 + (y')^2 } \] so let's add 1 to what you got: \[ \frac{x^6}{16} + x^{-6} - \frac{1}{2} +1 \] and simplify to \[ \frac{x^6}{16} + x^{-6} + \frac{1}{2} \] to make this easier to work with, multiply and divide by x^6 \[\frac{1}{x^6} \left( x^6 \left(\frac{x^6}{16} + x^{-6} + \frac{1}{2}\right) \right) \]

OpenStudy (phi):

distribute the x^6

OpenStudy (phi):

this will get rid of the x^-6 we will also factor out 1/16 from the inner expression, to get rid of the fractions

OpenStudy (phi):

you should get \[ \frac{1}{x^6} \left( \frac{x^{12}}{16} + 1 + \frac{x^6}{2} \right) \] if you also factor out 1/16 you get \[ \frac{1}{16x^6} \left( x^{12}+ 16 + 8x^6 \right) \] or \[ \frac{1}{16x^6} \left( x^{12}+ 8x^6+ 16 \right) \]

OpenStudy (phi):

notice that you have a perfect square. can you factor it ?

OpenStudy (anonymous):

Factor the x^12+8x^6+16?

OpenStudy (anonymous):

Is it (x^6+4)?

OpenStudy (phi):

yes (x^6+4)^2

OpenStudy (phi):

that means \[ \sqrt{1 + (y')^2 } = \sqrt{ \frac{(x^6+4)^2}{16x^6} }\] can you simplify that ?

OpenStudy (anonymous):

Um x^2+4/4x^6?

OpenStudy (anonymous):

Oops x^6+4/4x^6

OpenStudy (phi):

almost. the bottom x^6 should be x^3

OpenStudy (anonymous):

Because of the square root ?

OpenStudy (phi):

yes. \[ \sqrt{x^6} = x^3 \] and \( x^3 \cdot x^3 = x^6 \)

OpenStudy (phi):

or if you use \[ \sqrt{x^6} = \left(x^6\right)^{\frac{1}{2}} = x^{6\cdot\frac{1}{2}}= x^3 \]

OpenStudy (phi):

now you are ready to find the arc length \[ \int \sqrt{1 + (y')^2 } dx =\int_2^{3} \frac{(x^6+4)}{4x^3} \ dx\] we can simplify that to \[ \int_2^{3} \frac{x^3}{4}+x^{-3} \ dx\]

OpenStudy (anonymous):

So the integral is x^4/16-x^-4/4

OpenStudy (phi):

ok on x^4 / 16 but when you integrate you add +1 to the exponent. -3+1 is not -4

OpenStudy (anonymous):

Oh sorry x^-2

OpenStudy (phi):

so the integral is \[ \frac{x^4}{16} - \frac{1}{2} x^{-2} \]

OpenStudy (phi):

now evaluate it from x=2 to x= 3 to find the arc length.